Sarah is flying a kite. She is 5 ft. tall. The kite string makes a 38 degree angle of elevation with the horizontal and the kite is 99 ft. off the ground. How long is the kite string to the nearest tenth of a foot?
Do you subtract her height from the height of the kite from the ground to answer this question?

Answer by stanbon(75887)

She is 5 ft. tall. The kite string makes a 38 degree angle of elevation with the horizontal and the kite is 99 ft. off the ground. How long is the kite string to the nearest tenth of a foot?
Do you subtract her height from the height of the kite from the ground to answer this question?
---------
I would say yes.
----
Draw a picture of the situation.
You have a right triangle with
height = 99-5 = 94 ft
angle at her hand = 38 degrees
-------
Solve for the hypotenuse:
sin(38) = 94/hypotenuse
--------
hypotenuse = 94/sin(38) = 152.38 feet
----------------------
Cheers,
Stan H.
======================